# Density estimation via mixture discrepancy and moments

### v1

@@ Abstract
With the aim of generalizing histogram statistics to higher dimensional cases, density estimation via discrepancy based sequential partition (DSP) has been proposed [D. Li, K. Yang, W. Wong, Advances in Neural Information Processing Systems (2016) 1099–1107] to learn an adaptive piecewise constant approximation defined on a binary sequential partition of the underlying domain, where the star discrepancy is adopted to measure the uniformity of particle distribution. However, the calculation of the star discrepancy is NP-hard and it does not satisfy the reflection invariance and rotation invariance either. To this end, we use the mixture discrepancy and the comparison of moments as a replacement of the star discrepancy, leading to the density estimation via mixture discrepancy based sequential partition (DSP-mix) and density estimation via moments based sequential partition (MSP), respectively. Both DSP-mix and MSP are computationally tractable and exhibit the reflection and rotation invariance. Numerical experiments in reconstructing the $d$-D mixture of Gaussians and Betas with $d=2,3,\dots,6$ demonstrate that DSP-mix and MSP both run approximately ten times faster than DSP while maintaining the same accuracy.
@@
Density estimation; Adaptive sequential partition; Star discrepancy; Mixture discrepancy; Moment; Hausdorff moment problem; mixture of Gaussians; mixtures of Betas
@@ 1 Introduction
Density estimation is a classical problem in statistics and scientific computing. Given $N$ independent and identically distributed observations $\bm{x}_{1},\bm{x}_{2},\dots,\bm{x}_{N}$, which follow the probability density $p(\bm{x})$. This paper focuses on how to estimate $p(\bm{x})$ from $\bm{x}_{1},\bm{x}_{2},\dots,\bm{x}_{N}$ via a nonparametric and adaptive domain partition approach. In nonparametric models, a widely used method is the kernel density estimation with the magic lying in how to balance the bias and variance of the estimator [1, 2]. Another simple and classical approach is the histogram statistics, a uniform piecewise constant density estimator. Due to the curse of dimensionality, it is difficult to apply the histogram to high-dimensional problems [3]. An adaptive nonparametric density estimation via Discrepancy based Sequential Partition (DSP) was proposed to reconstruct an adaptive piecewise constant approximation of $p(\bm{x})$ [4] and DSP is expected to handle higher dimensions [5]. DSP uses the star discrepancy, a concept from the quasi-Monte Carlo method, to measure the uniformity of the observations [6, 7]. When the star discrepancy of the observations within a sub-domain is small, it can be reasonable to think that their distribution approaches the uniform distribution and thus the corresponding probability density can be regarded as approximately constant in the sub-domain.
For a point set $\left(\bm{y}_{1},\dots,\bm{y}_{n}\right)\subset\left[0,1\right]^{d}$, its star discrepancy reads
@@
and measures the uniformity of the points by using the infinity norm of the difference between the cumulative distribution function of the discrete point measure $\frac{1}{n}\sum_{i=1}^{n}\delta_{\bm{y}_{i}}$ and that of the uniform measure. Unfortunately, the star discrepancy possesses some drawbacks as follows [8].
  • 1.
    The star discrepancy is difficult to calculate: The calculation of the star discrepancy is NP-hard and requires the external approximate solver [9, 10]. When the sample size $N$ is large, the calculation is rather time-consuming (As shown in Tables 4 and 5, DSP needs dozens of seconds for $N=1\times 10^{6}$);
  • 2.
    The star discrepancy does not satisfy reflection invariance [8, 11]: The point sets in Figures 1(a) and 1(b) are symmetric about the line $x_{1}=0.5$, but the values of their star discrepancies are not equal, which is counterintuitive. The reason for this phenomenon lies in the fact that the origin $\boldsymbol{0}$ is distinctive in the definition of the star discrepancy;
  • 3.
(a) $D^{*}=\frac{23}{48}$, $D^{\text{mix}}=\frac{2719}{18432}$.
(b) $D^{*}=\frac{9}{16}$, $D^{\text{mix}}=\frac{2719}{18432}$.
Figure 1: Reflection invariance test, $d=2,N=3$: Figures 1(a) and 1(b) are symmetric about $x_{1}=0.5$. Intuitively, the reflection transformation should not change the uniformity of points. However, the star discrepancies of Figures 1(a) and 1(b) are different.
(a) $D^{*}=\frac{13}{24},D^{\text{mix}}=\frac{301}{1024}$.
(b) $D^{*}=\frac{7}{16},D^{\text{mix}}=\frac{301}{1024}$.
Figure 2: Rotation invariance test, $d=2,N=3$: Figure 2(b) is obtained by rotating Figure 2(a) 90 degrees clockwise. Intuitively, rotation should not change the uniformity of the points. However, the star discrepancies of Figure 2(a) and 2(b) are different.
In view of the drawbacks of the star discrepancy, this paper contributes two new uniformity measurement strategies for density estimation.
  • 1.
    One is the mixture discrepancy, another kind of discrepancy. For a point set $\left(\bm{y}_{1},\dots,\bm{y}_{n}\right)\subset\left[0,1\right]^{d}$, its mixture discrepancy has an analytical expression,
    $D^{\text{mix}}(\left(\bm{y}_{1},\ldots,\bm{y}_{n}\right))^{2}$ (2)
    $=$ $\left(\frac{19}{12}\right)^{d}-\frac{2}{n}\sum_{i=1}^{n}\prod_{j=1}^{d}\left(\frac{5}{3}-\frac{1}{4}\left|y_{ij}-\frac{1}{2}\right|-\frac{1}{4}\left|y_{ij}-\frac{1}{2}\right|^{2}\right)$
    $+\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{k=1}^{n}\prod_{j=1}^{d}\left(\frac{15}{8}-\frac{1}{4}\left|y_{ij}-\frac{1}{2}\right|-\frac{1}{4}\left|y_{kj}-\frac{1}{2}\right|-\frac{3}{4}\left|y_{ij}-y_{kj}\right|+\frac{1}{2}\left|y_{ij}-y_{kj}\right|^{2}\right),$
@@
    where $\bm{y}_{i}=(y_{i1},\dots,y_{id})$. It is thus easy to calculate, and possesses good theoretical properties. We use the same algorithm framework as DSP except for replacing the star discrepancy with the mixture discrepancy. The resulting method is dubbed the DSP-mix method;
  • 2.
    The other is the density estimation via moments based sequential partition (MSP), which still adopts the partitioning framework in DSP, but the way to measure the uniformity of the observations is changed to compare the gap between the moments of the particles and the moments of the uniform density.
@@
Whether the density function can be uniquely determined by the moments of all orders is the classical moment problem in probability theory [12, 13, 14]. For the case where the probability density is defined on a compact set, this problem is called the Hausdorff moment problem, and the answer is positive [15]. When the moments of all orders of the observations on the compact set are close to those of the uniform distribution, we can consider that the observations are approximately uniformly distributed. MSP is easy to calculate and satisfies the reflection and rotation invariance. Numerical experiments demonstrate that both DSP-mix and MSP achieve about ten times acceleration compared with DSP while maintaining the same accuracy.
@@
2 The DSP-mix method
DSP uses an adaptive piecewise constant density estimator to reconstruct the $d$-D probability density function from the observations [4]. DSP-mix inherits the algorithm framework of DSP and only differs in the measure of uniformity. Without loss of generality, assume that the observations $\bm{x}_{i}\in\Omega=\left[0,1\right]^{d},\,i=1,2,\dots,N$, and the computational domain $\Omega$ is decomposed into $L$ sub-domains,
@@ 2 The DSP-mix method
DSP-mix tries to approximate the density function $p(\bm{x})$ with a piecewise constant function,
$\hat{p}(\bm{x})=\sum_{l=1}^{L}c_{l}\mathbbm{1}_{\Omega_{l}}(\bm{x}),$ (4)
where
$c_{l}=\frac{1}{N|\Omega_{l}|}\sum_{i=1}^{N}\mathbbm{1}_{\Omega_{l}}(\bm{x}_{i}).$ (5)
To be more specific, DSP-mix adopts a sequential binary partitioning strategy to obtain the decomposition (3) during which the mixture discrepancy is adopted to measure the uniformity. When the mixture discrepancy of the observations within the sub-domain $\Omega_{l}$ is less than a predetermined threshold, it can be approximately considered that the observations within this sub-domain are uniformly distributed, and the corresponding density function can be approximated as a constant. Algorithm 1 presents the detailed process. It can be readily seen there that if there exists a sub-domain $\Omega_{l}$ fails to pass the uniformity test (See Lines 7-14), then $\Omega_{l}$ is split along a certain coordinate and two new sub-domains, $\Omega_{l}^{(1)}$ and $\Omega_{l}^{(2)}$, are generated. Subsequently, these two newly generated sub-domains are combined with the remaining sub-domains to update the sub-domain set $\mathcal{P}_{L+1}$ (see Line 17).
Algorithm 1 The DSP-mix method.
Input: The domain $\Omega$, particle set $\bm{x}_{i},\,i=1,2,\dots,N$, parameters $\theta,m$.
Output: An adaptive piecewise constant approximation of probability density function $p(\bm{x})$.
$L=1,\Omega_{1}=\Omega,\mathcal{P}_{1}=\left\{\Omega_{1}\right\}$;
@@
7:         if $S_{l}$ cannot be approximated as a uniform distribution then
@@
9:              Choose a splitting node $\widetilde{s}_{i_{0},j_{0}}^{(l)}$ according to Eq. (6);
10:              Divide $\Omega_{l}$ into $\Omega_{l}^{(1)}\cup\Omega_{l}^{(2)}$ as Eq. (7);
@@
22:Calculate piecewise constants $c_{l},\,l=1\dots,L$ according to Eq. (5) and partition $\mathcal{P}_{L}$;
@@
    Where to split: We adopt $m-1$ dividing points along each coordinate, and choose the uniformly distributed points for simplicity: $s_{i_{0},j_{0}}^{(l)}=a_{j_{0}}^{(l)}+\frac{i_{0}}{m}\left(b_{j_{0}}^{(l)}-a_{j_{0}}^{(l)}\right)$ with $i_{0}=1,2,\dots,m-1$ and $j_{0}=1,2,\dots,d$. Accordingly, DSP-mix splits $\Omega_{l}$ using the dividing point $\widetilde{s}_{i_{0},j_{0}}^{(l)}$ that divides $\Omega_{l}$ into $\Omega_{l}^{(1)}$ and $\Omega_{l}^{(2)}$. Here $\widetilde{s}_{i_{0},j_{0}}^{(l)}$ is selected according to
@@
    $\Omega_{l}^{(1)}=\prod_{j=1}^{j_{0}-1}\left[a_{j}^{(l)},b_{j}^{(l)}\right]\times\left[a_{j_{0}}^{(l)},s_{i_{0},j_{0}}^{(l)}\right]\times\prod_{j=j_{0}+1}^{d}\left[a_{j}^{(l)},b_{j}^{(l)}\right],$ (7)
    $\Omega_{l}^{(2)}=\Omega_{l}\backslash\Omega_{l}^{(1)}.$
@@
    Whether to split: The linear scaling $\Omega_{l}\rightarrow\left[0,1\right]^{d}$ is performed first before using the discrepancy as the uniformity measurement. We scale $S_{l}$ to $\widetilde{S}_{l}=\{(\frac{y^{(l)}_{i1}-a_{1}^{(l)}}{b_{1}^{(l)}},\dots,\frac{y^{(l)}_{id}-a_{d}^{(l)}}{b_{d}^{(l)}})\}_{i=1}^{n}$. The partitioning terminates when the mixture discrepancy of particles that fall within $\Omega_{l}$ satisfying
    $D^{\text{mix}}(\widetilde{S}_{l})\leq\frac{\theta\sqrt{N}}{n}.$ (8)
    DSP-mix uses the mixture discrepancy $D^{\text{mix}}(\widetilde{S}_{l})$ to measure the uniformity of $S_{l}$ in Line 6 of Algorithm 1, while DSP uses the star discrepancy $D^{*}(\widetilde{S}_{l})$ instead. In Line 7 of Algorithm 1, DSP-mix considers $S_{l}$ inapproximable as a uniform distribution when $D^{\text{mix}}(\widetilde{S}_{l})>\frac{\theta\sqrt{N}}{n}$, while DSP holds the same for $D^{*}(\widetilde{S}_{l})>\frac{\theta\sqrt{N}}{n}$. The parameter $\theta$ governs the depth of partition. The smaller $\theta$ is, the finer the partitioning it induces.

### v2

@@ Abstract
With the aim of generalizing histogram statistics to higher dimensional cases, density estimation via discrepancy based sequential partition (DSP) has been proposed to learn an adaptive piecewise constant approximation defined on a binary sequential partition of the underlying domain, where the star discrepancy is adopted to measure the uniformity of particle distribution. However, the calculation of the star discrepancy is NP-hard and it does not satisfy the reflection invariance and rotation invariance either. To this end, we use the mixture discrepancy and the comparison of moments as a replacement of the star discrepancy, leading to the density estimation via mixture discrepancy based sequential partition (DSP-mix) and density estimation via moment-based sequential partition (MSP), respectively. Both DSP-mix and MSP are computationally tractable and exhibit the reflection and rotation invariance. Numerical experiments in reconstructing Beta mixtures, Gaussian mixtures and heavy-tailed Cauchy mixtures up to 30 dimension are conducted, demonstrating that MSP can maintain the same accuracy compared with DSP, while gaining an increase in speed by a factor of two to twenty for large sample size, and DSP-mix can achieve satisfactory accuracy and boost the efficiency in low-dimensional tests ($d\leq 6$), but might lose accuracy in high-dimensional problems due to a reduction in partition level.
@@
Density estimation; Adaptive sequential partition; Discrepancy; Hausdorff moment problem; Kullback-Leibler divergence; Hellinger distance; Beta mixtures; Gaussian mixtures; Cauchy mixtures
@@ 1 Introduction
As a fundamental problem in statistics, density estimation aims at constructing an estimate of their common density function $q(\bm{y})$ for given $N$ observations $\bm{y}_{1},\bm{y}_{2},\dots,\bm{y}_{N}$. It has various applications in uncertainty quantification [1], Bayesian inference [2] and scientific computing [3, 4, 5, 6]. A simple and classical approach is the histogram statistics as a uniform piecewise constant density estimator. Unfortunately, it is rather difficult to apply the histogram to high-dimensional problems as both bin size and required sample size grow exponentially in dimension, which is well known as the curse of dimensionality (CoD) [7, 3].
This paper focuses on nonparametric approaches for multivariate density estimation, especially for the case when the dimension $d$ is moderately large, say, 6306–30 [8, 9]. The benefit of nonparametric methods is their ability to achieve estimation optimality for any input distribution as more data are observed. A widely used method is the kernel density estimation (KDE), treating each data point as the center of a kernel function like the Gaussian functions and smoothed splines [7]. But the accuracy of the kernel estimator becomes very sensitive to the choice of the window size and the shape of the kernel [10], and the magic lies in how to balance the bias and variance of the estimator, like using the smoothing properties of linear diffusion processes [11] or choosing adaptive bandwidths [12, 10]. Another approach is based on tensor decompositions, with huge advantages in that the computational cost of the construction, the storage requirements and the operations required for conditional distribution method sampling from the distributional approximation all scale linearly with dimension [4, 5, 13]. However, it still relies heavily on the low-rank assumption of the underlying density, whereas the detection and characterization of local features of multivariate density estimation are usually prohibitive [8]. Machine learning approaches have recently been drawing a growing attention, e.g., the density estimation through normalizing flows [14] and deep generative neural networks [15]. In practice, they require extensive tuning to perform well [16].
Here we mainly discuss a data-driven density estimation method, termed tree-based density estimation. This features a class of estimators which employs simple and flexible binary partitions to adapt to the underlying density function, along with the decision tree using stopping times in a data-driven way [2]. The decision tree based method has a great potential to overcome the burden imposed by the high dimensionality, especially when the dimension is only moderately large and the density function exhibits certain spatial features that can be leveraged of [16]. Essentially, it is equivalent to cluster the samples into small nonintersecting sets, each supported by a tree-structured density [17]. An adaptive nonparametric density estimation can be constructed by either the Bayesian Sequential Partition (BSP) [8, 16] or the Discrepancy based Sequential Partition (DSP) [18]. Compared with BSP, the computational cost of DSP is much cheaper due to its greedy construction, and is fully capable to handle higher dimension [6].
To be more specific, DSP uses the star discrepancy, a concept from the quasi-Monte Carlo method, to measure the uniformity of the observations [19, 20]. For a point set $\left(\bm{y}_{1},\dots,\bm{y}_{n}\right)\subset\left[0,1\right]^{d}$, its star discrepancy reads
@@
and characterizes the uniformity of the points by using the infinity norm of the difference between the cumulative distribution function of the discrete point measure $\frac{1}{n}\sum_{i=1}^{n}\delta_{\bm{y}_{i}}$ and that of the uniform measure. When the star discrepancy of the observations within a sub-domain is small, it can be reasonable to think that their distribution approaches the uniform distribution and thus the corresponding probability density can be regarded as approximately constant in the sub-domain.
The major bottleneck of DSP lies in the complexity of calculating the star discrepancy of pointsets in the $d$-dimensional unit cube, which has been proved to be an NP-hard problem [21]. As a consequence, the external approximate solver [22, 23, 24] is mandatory, but is rather time-consuming when the sample size $N$ is large. Besides, the star discrepancy lacks both reflection and rotation invariance [20, 25, 26] since the origin $\boldsymbol{0}$ is distinctive in the definition (1) of the star discrepancy. In view of such drawbacks, this paper contributes two new uniformity measurement strategies for density estimation, both of which seek a relaxation of DSP. One is to adopt the mixture discrepancy, another kind of discrepancy [20, 25, 26]. For a point set $\left(\bm{y}_{1},\dots,\bm{y}_{n}\right)\subset\left[0,1\right]^{d}$, its mixture discrepancy has an analytical expression,
$\left(D^{\text{mix}}(\mathbf{y}_{1},\ldots,\mathbf{y}_{n})\right)^{2}=$ $\left(\frac{19}{12}\right)^{d}-\frac{2}{n}\sum_{i=1}^{n}\prod_{j=1}^{d}\left(\frac{5}{3}-\frac{1}{4}\left|y_{ij}-\frac{1}{2}\right|-\frac{1}{4}\left|y_{ij}-\frac{1}{2}\right|^{2}\right)$ (2)
$+\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{k=1}^{n}\prod_{j=1}^{d}\left(\frac{15}{8}-\frac{1}{4}\left|y_{ij}-\frac{1}{2}\right|-\frac{1}{4}\left|y_{kj}-\frac{1}{2}\right|\right$.
$\left.-\frac{3}{4}\left|y_{ij}-y_{kj}\right|+\frac{1}{2}\left|y_{ij}-y_{kj}\right|^{2}\right),$
@@
where $\bm{y}_{i}=(y_{i1},\dots,y_{id})$. It is thus easy to calculate, and possesses good theoretical properties. The interested readers are referred to [20, 25, 26] for more details about the mixture discrepancy including the motivation and the calculation. We use the same algorithm framework as DSP except for replacing the star discrepancy with the mixture discrepancy. The resulting method is dubbed the DSP-mix method.
@@
The other is the density estimation via moment-based sequential partition (MSP), which still adopts the partitioning framework in DSP, but the way to measure the uniformity of the observations is replaced by comparing the gap in moments between the particle distribution and the uniform density. Whether the density function can be uniquely determined by the moments of all orders is the classical moment problem in probability theory [27, 28, 29]. For the case where the probability density is defined on a compact set, this problem is called the Hausdorff moment problem, and the answer is positive [30]. When the moments of truncated orders of the observations on the compact set are close to those of the uniform distribution, we can consider that the observations are approximately uniformly distributed. Numerical experiments in reconstructing Beta mixtures, Gaussian mixtures and heavy-tailed Cauchy mixtures up to 30 dimension are conducted, demonstrating that MSP can maintain the same accuracy compared with DSP, while gaining an increase in speed by a factor of two to twenty for large sample size, and DSP-mix can achieve satisfactory accuracy in low-dimensional tests ($d\leq 6$) and run significantly faster than DSP, but might lose accuracy in high-dimensional problems due to a reduction in partition level.
@@
2 Discrepancy sequential partition via star discrepancy or mixed discrepancy
Without loss of generality, assume that the observations $\bm{y}_{i}\in\Omega,\,i=1,2,\dots,N$, and the computational domain $\Omega$ is decomposed into $L$ non-intersecting sub-domains,
@@ 2 Discrepancy sequential partition via star discrepancy or mixed discrepancy
DSP uses an adaptive piecewise constant density estimator to reconstruct the $d$-D probability density function from the observations [18],
$\hat{p}(\bm{y})=\sum_{l=1}^{L}c_{l}\mathbbm{1}_{\Omega_{l}}(\bm{y}),\quad c_{l}=\frac{1}{N|\Omega_{l}|}\sum_{i=1}^{N}\mathbbm{1}_{\Omega_{l}}(\bm{y}_{i}).$ (4)
As presented in Figure 1, DSP adopts a sequential binary partitioning strategy to obtain the decomposition (3).
Figure 1: An illustration of tree-based density estimation.
2.1 Setting of tree-based partitioning
Algorithm 1 The framework of tree-based density estimation.
Input: The domain $\Omega$, particle set $\bm{y}_{i},\,i=1,2,\dots,N$, parameters $\theta,m$.
Output: An adaptive piecewise constant approximation of probability density function $q(\bm{y})$.
[…]
@@ 2.1 Setting of tree-based partitioning
7:   if $S_{l}$ cannot be approximated as a uniform distribution, e.g., $D^{\ast}(\widetilde{S}_{l})>\frac{\theta\sqrt{N}}{n_{l}}$ then.
@@
9:     Choose a splitting node $\widetilde{s}_{i_{0},j_{0}}^{(l)}$ according to
$\widetilde{s}_{i_{0},j_{0}}^{(l)}=\underset{s_{i_{0},j_{0}}^{(l)}}{\arg\max}\left|\frac{n_{i_{0},j_{0}}}{n}-\frac{i_{0}}{m}\right|$
10:     Divide $\Omega_{l}$ into $\Omega_{l}^{(1)}\cup\Omega_{l}^{(2)}$ as
$\Omega_{l}^{(1)}=\prod_{j=1}^{j_{0}-1}\left[a_{j}^{(l)},b_{j}^{(l)}\right]\times\left[a_{j_{0}}^{(l)},s_{i_{0},j_{0}}^{(l)}\right]\times\prod_{j=j_{0}+1}^{d}\left[a_{j}^{(l)},b_{j}^{(l)}\right],\quad\Omega_{l}^{(2)}=\Omega_{l}\backslash\Omega_{l}^{(1)}$
@@
22:Calculate a piecewise constant density estimator $\hat{p}(\bm{y})$ corresponding to the partition $\mathcal{P}_{L}$ by
$\hat{p}(\bm{y})=\sum_{l=1}^{L}c_{l}\mathbbm{1}_{\Omega_{l}}(\bm{y}),\quad c_{l}=\frac{1}{N|\Omega_{l}|}\sum_{i=1}^{N}\mathbbm{1}_{\Omega_{l}}(\bm{y}_{i})$
When the star discrepancy given in Eq. (1) of the observations within the sub-domain $\Omega_{l}$ is less than a predetermined threshold, it can be approximately considered that the observations within this sub-domain are uniformly distributed, and the corresponding density function can be approximated as a constant. Algorithm 1 presents the detailed process. It can be readily seen there that if there exists a sub-domain $\Omega_{l}$ that fails to pass the uniformity test (see Lines 7-14), then $\Omega_{l}$ is split along a certain coordinate and two new sub-domains, $\Omega_{l}^{(1)}$ and $\Omega_{l}^{(2)}$, are generated. Subsequently, these two newly generated sub-domains are combined with the remaining sub-domains to update the sub-domain set $\mathcal{P}_{L+1}$ (see Line 17).
@@
    Where to split: We adopt m1m-1 dividing points along each coordinate, and choose the uniformly distributed points for simplicity: $s_{i_{0},j_{0}}^{(l)}=a_{j_{0}}^{(l)}+\frac{i_{0}}{m}\left(b_{j_{0}}^{(l)}-a_{j_{0}}^{(l)}\right)$ with $i_{0}=1,2,\dots,m-1$ and $j_{0}=1,2,\dots,d$. Accordingly, DSP splits $\Omega_{l}$ using the dividing point $\widetilde{s}_{i_{0},j_{0}}^{(l)}$ that divides $\Omega_{l}$ into $\Omega_{l}^{(1)}$ and $\Omega_{l}^{(2)}$. Here $\widetilde{s}_{i_{0},j_{0}}^{(l)}$ is selected according to
@@
    $\Omega_{l}^{(1)}=\prod_{j=1}^{j_{0}-1}\left[a_{j}^{(l)},b_{j}^{(l)}\right]\times\left[a_{j_{0}}^{(l)},s_{i_{0},j_{0}}^{(l)}\right]\times\prod_{j=j_{0}+1}^{d}\left[a_{j}^{(l)},b_{j}^{(l)}\right],\quad\Omega_{l}^{(2)}=\Omega_{l}\backslash\Omega_{l}^{(1)}.$ (6)
@@
    Whether to split: The linear scaling $\Omega_{l}\rightarrow\left[0,1\right]^{d}$ is performed first before using the discrepancy as the uniformity measurement. We scale $S_{l}$ to $\widetilde{S}_{l}=\{(\frac{y^{(l)}_{i1}-a_{1}^{(l)}}{b_{1}^{(l)}},\dots,\frac{y^{(l)}_{id}-a_{d}^{(l)}}{b_{d}^{(l)}})\}_{i=1}^{n}$. The partitioning terminates when the discrepancy of particles that fall within $\Omega_{l}$ satisfying
    $D^{\ast}(\widetilde{S}_{l})\leq\frac{\theta\sqrt{N}}{n_{l}}.$ (7)
    The parameter $\theta$ governs the depth of partition. The smaller $\theta$ is, the finer the partitioning it induces.
2.2 Theoretical analysis of DSP
We can use the $\ast$-total variation distance to measure the difference between the density estimator $\hat{p}(\bm{y})$ and the empirical measure $\mu=\frac{1}{N}\sum_{i=1}^{N}\delta_{\bm{y}_{i}}$. For two measures $\mu$ and $\nu$, the $\ast$-total variation distance is defined as
[…]
where $\mu(A)=\int_{\Omega}1_{A}(\bm{y})\textup{d}\mu(\bm{y})$. A rough error estimator is given as follows.
Theorem 2.1.
For the partition $\Omega=\cup_{l=1}^{L}\Omega_{l}$, suppose $D^{*}(\widetilde{S}_{l})\leq\theta\sqrt{N}/n_{l}$ in each subregion $\Omega_{l}$, then it has
$D^{\ast}[\mu,\hat{p}]\leq\frac{L\theta}{\sqrt{N}}.$ (8)
Thus for any function $f$ with bounded variation $V_{HK}(f;\Omega)$ in the sense of Hardy and Krause, the transport cost has an upper bound
$\Big|\int_{\Omega}f\hat{p}\textup{d}\bm{y}-\int_{\Omega}f\textup{d}\mu\Big|\leq\frac{L\theta}{\sqrt{N}}V_{HK}(f;\Omega).$ (9)
Proof.
Denote by $\{\bm{y}_{i}\in A\}$ the count of particles in $A$, and $\bm{y}_{i}^{(l)}$ the particles in the subregion $\Omega_{l}$. For arbitrary $A\subset\Omega$, it has
$\mu(A)=\frac{1}{N}\#\{\bm{y}_{i}\in A\},\quad\hat{p}(A)=\sum\limits_{l=1}^{L}\frac{n_{l}}{N}\frac{\textup{vol}(A\cap\Omega_{l})}{\textup{vol}(\Omega_{l})}.$
Since $A\cap\Omega_{l}$ is still anchored at the left corner of $\Omega_{l}$, it yields that
$\begin{split}|\mu(A)-\nu(A)|&=\left|\frac{1}{N}\#\{\bm{y}_{i}\in A\}-\sum_{l=1}^{L}\frac{n_{l}}{N}\frac{\textup{vol}(A\cap\Omega_{l})}{\textup{vol}(\Omega_{l})}\right|\\ &=\left|\frac{1}{N}\sum_{l=1}^{L}\#\{\bm{y}_{i}^{(l)}\in A\cap\Omega_{l}\}-\sum_{l=1}^{L}\frac{n_{l}}{N}\frac{\textup{vol}(A\cap\Omega_{l})}{\textup{vol}(\Omega_{l})}\right|\\ &\leq\frac{1}{N}\sum_{l=1}^{L}n_{l}\left|\frac{\#\{\bm{y}_{i}^{(l)}\in A\cap\Omega_{l}\}}{n_{l}}-\frac{\textup{vol}(A\cap\Omega_{l})}{\textup{vol}(\Omega_{l})}\right|\\ &\leq\frac{1}{N}\sum_{l=1}^{L}n_{l}D^{*}(\widetilde{S}_{l})\leq\frac{1}{N}\sum_{l=1}^{L}n_{l}\frac{\theta\sqrt{N}}{n_{l}}=\frac{L\theta}{\sqrt{N}}.\end{split}$
By taking supremum over $A\subset\Omega$, it arrives at $D^{*}[\mu,\hat{p}]\leq\frac{L\theta}{\sqrt{N}}$. Eq. (9) is a direct consequence of the Koksma-Hlawka inequality. ∎
Since the partition level $L$ actually depends on $\theta$ and $N$, the above rough bound is useful only when $L$ is not so large. In particular, the total variation $V_{HK}(f;\Omega)$ scales linearly with $d$. For $f(\bm{y})=\sum_{i=1}^{d}y_{i}^{2}$ or $f(\bm{y})=\sum_{j\neq i}\sum_{i=1}^{d}y_{i}y_{j}$, $V_{HK}(f;\Omega)$ scales quadratically with $d$. As a result, Eq. (9) controls the discrepancy in the first and second moments of $\hat{p}$ and $\mu$.
2.3 Relaxation by the mixed discrepancy
Despite that there are some efficient heuristic algorithms for calculating the star discrepancy, like the TA-improved method [23] and the ODE annealing approach [24], the computational cost dramatically grows as sample size $N$ and dimension $d$ increase. Apart from the huge computational complexity, the star discrepancy possesses two drawbacks as follows [25].
(a) $D^{*}=\frac{23}{48}$, $D^{\text{mix}}=\frac{2719}{18432}$.
(b) $D^{*}=\frac{9}{16}$, $D^{\text{mix}}=\frac{2719}{18432}$.
Figure 2: Reflection invariance test, $d=2,N=3$: Figures 2(a) and 2(b) are symmetric about $x_{1}=0.5$. Intuitively, the reflection transformation should not change the uniformity of points. However, the star discrepancies of Figures 2(a) and 2(b) are different.
(a) $D^{*}=\frac{13}{24},D^{\text{mix}}=\frac{301}{1024}$.
(b) $D^{*}=\frac{7}{16},D^{\text{mix}}=\frac{301}{1024}$.
Figure 3: Rotation invariance test, $d=2,N=3$: Figure 3(b) is obtained by rotating Figure 3(a) 90 degrees clockwise. Intuitively, rotation should not change the uniformity of the points. However, the star discrepancies of Figure 3(a) and 3(b) are different.
  1. (1)
    The star discrepancy does not satisfy reflection invariance [25, 26]: The point sets in Figures 2(a) and 2(b) are symmetric about the line $x_{1}=0.5$, but the values of their star discrepancies are not equal, which is counterintuitive. The reason for this phenomenon lies in the fact that the origin $\boldsymbol{0}$ is distinctive in the definition of the star discrepancy;
  2. (2)
To fulfill reflection and rotation invariances, we adopt the mixed discrepancy (2) as introduced in [20, 25, 26]. The resulting DSP-mix inherits the algorithm framework of DSP and only differs in the measure of uniformity, where the mixture discrepancy $D^{\text{mix}}(\widetilde{S}_{l})$ to adopted to measure the uniformity of $S_{l}$ in Line 6 of Algorithm 1,
$D^{\text{mix}}(\widetilde{S}_{l})\leq\frac{\theta\sqrt{N}}{n_{l}}.$ (10)